# Generative World Explorer

### v1

@@ Abstract
Planning with partial observation is a central challenge in embodied AI. A majority of prior works have tackled this challenge by developing agents that physically explore their environment to update their beliefs about the world state. In contrast, humans can imagine unseen parts of the world through a mental exploration and revise their beliefs with imagined observations. Such updated beliefs can allow them to make more informed decisions, without necessitating the physical exploration of the world at all times. To achieve this human-like ability, we introduce the Generative World Explorer (Genex), an egocentric world exploration framework that allows an agent to mentally explore a large-scale 3D world (e.g., urban scenes) and acquire imagined observations to update its belief. This updated belief will then help the agent to make a more informed decision at the current step. To train Genex, we create a synthetic urban scene dataset, Genex-DB. Our experimental results demonstrate that (1) Genex can generate high-quality and consistent observations during long-horizon exploration of a large virtual physical world and (2) the beliefs updated with the generated observations can inform an existing decision-making model (e.g., an LLM agent) to make better plans.
[Uncaptioned image] Website https://generative-world-explorer.github.io/
[Uncaptioned image] Code https://github.com/Beckschen/genex
Figure 1: We propose the Generative World Explorer Genex that allows an agent to imaginatively explore a large-scale 3D virtual world and acquire imagined observation to update its belief. In this example, agent 1 (sliver car) imaginatively explores to the perspective at the location of agent 2 (taxi) with Genex. The imagined observation at agent 2’s position (including the explored ambulance) revises agent 1’s belief and enables more informed decision making.
@@ 1 Introduction
Furthermore, the proposed Genex can be applied to the embodied decision making. With Genex, the agent is able to imagine hidden views via imaginatively exploration, and revise its belief. The revised belief allows the agent to take more informed actions. Technically, we define the agent’s behavior as an extension of POMDP with imagination-driven belief revision. Notably, the proposed Genex can naturally be extended to multi-agent scenarios, where one agent can mentally navigate to the positions of other agents and update its own beliefs based on imagined beliefs of the other agents.
@@ 2 Related Works
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) have proven effective in image generation. To render high-resolution images, the latent diffusion models (LDMs) (Rombach et al., 2022) are proposed to denoise in the latent space. Similarly, video diffusion models (Blattmann et al., 2023b; Wang et al., 2023a; Blattmann et al., 2023a) use VAE models to encode video frames and denoise in the latent space. For controllable synthesis, the conditional denoising autoencoder are implemented with text (Rombach et al., 2022; OpenAI, 2024) and various conditioning controls (Zhang et al., 2023; Sudhakar et al., 2024). We focus on video generation conditioned on the egocentric panoramic view of agent, as panorama (Li & Bansal, 2023; 2024) ensures coherence in the generated world, and the use of egocentric vision is de facto choice in many embodied tasks (Das et al., 2018; Sermanet et al., 2024).
Generative vision for embodied decision making. Decision-making in the physical world (Das et al., 2018; Sermanet et al., 2024) is a fundamental AI challenge. LLMs provide linguistic reasoning that aids decision-making (Hao et al., 2023; Min et al., 2024) and vision-language planning (Cen et al., 2024). World models offer predictive representations of future states to inform decisions, though early attempts (Ha & Schmidhuber, 2018; LeCun, 2022) focus on simple game agents and often lack commonsense reasoning about the physical world. Generative vision (OpenAI, 2024; Kondratyuk et al., 2024) and in-context learning (Bai et al., 2024; Zhang et al., 2024) offer new avenues for using video generation to guide real-world decision-making (Yang et al., 2024c). Several works focus on specific application domains such as autonomous driving (Hu et al., 2023; Wang et al., 2023b; 2024c) which limit their generality. Others like video in-context learning (Zhang et al., 2024) requires a known demonstration video, which is inefficient for decision-making. Action-conditioned video generation models (Du et al., 2024a; Yang et al., 2024b; c; Wang et al., 2024a; Bu et al., 2024; Souček et al., 2024; Du et al., 2024b) can directly synthesize visual plans for decision-making. These models, however, focus on state transition probabilities without explicitly modeling agent beliefs, which are crucial for reasoning about other objects/agents in partially observable environments.
@@ 3 Generative World Exploration
Figure 2: Genex is able to explore an imaginative world by generating imagined video outputs, given RGB observations, exploration direction, and distance as inputs (a). Genex, grounded in physical environment, can perform GPT-assisted imaginative exploration (b) and target-driven imaginative navigation (c).
@@ 3.1 Macro-Design of Genex
(a) Overview. As shown in  Fig. 2, the Genex framework enables agents to explore within an imaginative world by streaming video generation, based on current RGB observations and given exploration configurations. The RGB observation is represented as a panorama image sampled from any location in the world. A large multimodal model (LMM) serves as the pilot to set up exploration configurations, including any $360^{\circ}$ navigation direction and distance. Genex processes the input in two steps. First, it takes the exploration orientation to update the panorama forward view. Secondly, its built-in diffuser generates the forward navigation video. Both view update and diffuser are detailed in § 3.2.
@@ 3.2 Micro-Design of Genex
(a) Diffuser backbone. To support exploration illustrated in Fig. 2, we propose a video diffuser that can be seamlessly adapted to be a world explorer. Given an initial panorama image $x^{0}$ with camera position $p^{0}$, our objective is to generate a sequence of images $\{x^{1},\dots,x^{n}\}$ corresponding to a sequence of camera positions $\{p^{1},\dots,p^{n}\}$. The camera positions progress steadily forward, representing navigation in the world. Since the panorama image represents a 360-degree view, the generation should persist the information stored in previous frames to maintain world consistency throughout the sequence. Our model uses the pretrained stable video diffusion (SVD) (Blattmann et al., 2023a). The Transformer UNet (Ronneberger et al., 2015; Chen et al., 2021) architecture is as described in Blattmann et al. (2023b), where temporal convolution and attention layers are inserted after every spatial convolution and attention layer. The pipeline of our model is shown in Fig. 3 (a). Given an image condition $c$ (encoded from image $x^{0}$ using CLIP image Transformer (Radford et al., 2021)), the video diffusion algorithms learn a network $\epsilon_{\theta}$ to predict the noise added to the noisy image latent $z_{t}$ with $\mathcal{L}_{\text{noise}}=\|\epsilon_{\theta}(z_{t},c)-\epsilon_{t}\|^{2}$.
(b) Input image representation. Panorama images are optimal for generative exploration as they captures all perspectives from an egocentric viewpoint into a 2D image. Essentially, it represents a spherical polar coordinate system $\mathcal{S}$ on a 2D grid in the Cartesian coordinate system $\mathcal{P}$, as shown in Fig. 3 (b). The panorama image effectively stores every perspective of the world from a single location which preserves the global context during spatial navigation. This allows us to maintain consistency in world information from the conditional image, ensuring that the generated content aligns coherently with the surrounding environment. The panorama image also allows for rotational transformations, which facilitate world navigation by enabling us to rotate the image to face a different angle while preserving its original information. The rotation can be performed using Eq. 1:
@@
where $u$ and $v$ are positions on the 2D image plane, and $\phi$ and $\theta$ represent longitude and latitude in polar coordinates. The rotation function $\mathcal{R}$ applies a rotation to the spherical representation in any direction, simulating turning around during navigation. Additionally, a panorama image can be converted into a cubemap of six separate regular images, each representing a face of a cube (front, back, left, right, top, and bottom). This panorama-to-cube transformation enhances visual understanding by multimodal LLM agents. Full mathematical details of the equirectangular projection are in § A.1.
@@
$\mathcal{L}=\lambda\underbrace{||\mathcal{E}(\mathcal{T}(\mathcal{D}(z_{t}-\epsilon_{\theta}(z_{t},c))))-\mathcal{E}(\mathcal{T}(x_{0}))||^{2}}_{\mathcal{L}_{scl}}+\underbrace{(1-\lambda)||\epsilon_{\theta}(z_{t},c)-\epsilon_{t}||^{2}}_{\mathcal{L}_{\text{noise}}},$ (2)
@@ 4.1 Imagination-driven Belief Revision
Embodied agents operate under a POMDP framework (Puterman, 1994; Kaelbling et al., 1998; Nayak et al., 2024) . At each time step $t$, the agent’s world state (which represents the complete environment at this specific moment), $s^{t}\in S$, and action $a^{t}\in A$ determine the next world state via the transition probability ${T}(s^{t+1}|s^{t},a^{t})$. The agent’s given goal $g\in G$ (e.g., crossing the street) influences the reward $r^{t}=R(s^{t},a^{t},g)$, which drives the agent to achieve its objective. The agent receives an observation $o^{t}\in\Omega$ based on the observation model $O(o|s^{t})$ and maintains a belief, represented by a distribution $b(s)$, which is the agent’s internal estimate of the true state of the world. Its belief is updated with new observations, following the POMDP framework in Eq. 3:
$b^{t+M}(s^{t+M})=\prod_{t}^{M}\bigg{(}\underbrace{O(o^{t+1}|s^{t+1},a^{t})\sum_{s^{t}}T(s^{t+1}|s^{t},a^{t})}_{\text{Physical Exploration}}\bigg{)}b^{t}(s^{t})$ (3)
The decision $a^{t}$ made at any time $t$ becomes more informed as the agent gains a clearer understanding of its surroundings. By navigating through physical space, the agent gathers additional information about its environment, enabling more accurate assessments and better choices moving forward. However, physically traversing the space is inefficient, expensive, and even impossible in dangerous scenario. To streamline this process, we can use imagination as a medium for the agent to simulate outcomes without physically traversing. The key question becomes:
@@
Imagination-driven belief revision. We propose imagination-driven belief revision that uses imaginative exploration to enhance POMDP agents with a instant belief revision between time steps. In imagination, we freeze the time and create an imagined world, thus dropping the time variable $t$ and defining an imagination space with hat $\hat{}$ on the variables. Similar to physical navigation, the agent can make a sequence of imaginative actions $\hat{\mathbf{a}}=\{\hat{a}_{i}\in\hat{A}\}$ over imagination time step $I=\{1,...i,...,n\}$. With similar reward $\hat{R}$ encouraging agents to make sequential speculation on the unobserved world based on its initial belief and toward ultimate goal, it imagine novel observations in a previous unobserved world with $p_{\hat{\theta}}(\hat{o}^{i+1}|\hat{o}^{i},{\hat{a}^{i}})$. As a result, it can update its belief with Eq. 4:
$\hat{b}^{t}(s^{t})=\prod_{i}^{I}\bigg{(}\underbrace{p_{\theta}(\hat{o}^{i+1}|o^{i},\hat{a}^{i})}_{\text{Imaginative Exploration}}\bigg{)}b^{t}(s^{t})$ (4)
@@
The agent will make a action based on its belief and goal, with policy model $\pi(a^{t}|b^{t}(s^{t}),g)$. Through the revised belief, the agent is capable of making more informed decision toward $a^{*}$ with a more refined belief toward $b^{*}$, with more information on the true state of its surrounding environment.
In our work, we apply Genex for imaginative exploration and a LMM for the policy model $\pi$ and belief updater $b(s)$, mapping from observation to belief, with examples in Fig. 5 and system pipeline in § B.0.3.
@@ 4.2 Generalized to Multi-agent
Thus, we can adjust agent-$1$’s beliefs by aggregating the imagined belief counterpart for other $K-1$ agents.
@@ 5.1 Dataset construction
Genex-DB. We synthesize a large-scale dataset generated using Unity, Blender, and Unreal Engine. The full details are in § A.3. We create four distinct scenes, each representing a different visual style (Realistic, Animated, Low-Texture, and Geometry), shown in Fig. 6: We train a model with each dataset, and for the four resulting navigational video diffusers, we conduct cross-validation across all scenes to evaluate their generalization capabilities (detailed in § B.2).
@@
Genex-EQA. Through the proposed Genex model, the agents perform exploration autonomously, capable of tackling embodied tasks such as single agent scene understanding and multi-agent intersectional reasoning. Following the four conditions in § 4.3, we design over 200 scenarios in virtual physical engine to test various multimodal LLM agents embodied decision-making. We provide comprehensive details in § A.4.1. The dataset generally represent two scenarios:
@@ 5.2 Evaluation on Generation Quality
We adopt four metrics to evaluate the quality of video generation, including FVD (Unterthiner et al., 2019), SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018), and PSNR (Horé & Ziou, 2010). Evaluation details are in Appendix B.
@@
As a strong baseline, we develop a six-view navigator by training six separate diffusion model for each face of the cube, representing still a 360 view, independently (See Fig. 3 (b) six-view). The implementation detail is shown in § B.1. This baseline may align well with 2D diffusion models but stands in contrast to the panoramic approach, which is particularly effective at maintaining consistent environmental context. To enable a fair comparison with Genex in video quality evaluation, the six-view baseline predictions are reprojected into panoramas. As a result, Table 1 shows that our method achieves high generation quality and surpass six-view baseline in all metrics.
Figure 7: Example randomly sampled trajectory for cycle consistency. It forms a closed loop within the scene, with 9 rotations and in 15 meters distance.

### v3

# GenEx: Generating an Explorable World

@@ Abstract
Understanding, navigating, and exploring the 3D physical real world has long been a central challenge in the development of artificial intelligence. In this work, we take a step toward this goal by introducing GenEx, a system capable of planning complex embodied world exploration, guided by its generative imagination that forms expectations about the surrounding environments. GenEx generates high-quality, continuous 360-degree virtual environments, achieving high loop consistency and active 3D mapping over extended trajectories. Leveraging generative imagination, GPT-assisted agents can undertake complex embodied tasks, including goal-agnostic exploration and goal-driven navigation. Agents utilize imagined observations to update their beliefs, simulate potential outcomes, and enhance their decision-making. Training on the synthetic urban dataset GenEx-DB and evaluation on GenEx-EQA demonstrate that our approach significantly improves agents’ planning capabilities, providing a transformative platform toward intelligent, imaginative embodied exploration.
[Uncaptioned image] Website https://www.GenEx.world/
[Uncaptioned image] Code https://github.com/GenEx-world/genex
[Uncaptioned image] ArXiv https://arxiv.org/abs/2412.09624
Figure 1: GenEx explores an imaginative world, created from a single RGB image and brought to life as a generated video (top). With interactive and diverse generation, GenEx enables an agent to imaginatively explore a large-scale 3D world and acquire imagined observation to augment embodied intelligence (bottom).
@@ 1 Introduction
Furthermore, the proposed GenEx can be applied to the embodied decision making. With GenEx, the agent is able to imagine hidden views via imaginative exploration, and revise its belief. The revised belief allows the agent to take more informed actions. Technically, we define the agent’s behavior as an extension of POMDP with imagination-driven belief revision. Notably, the proposed GenEx can naturally be extended to multi-agent scenarios, where one agent can mentally navigate to the positions of other agents and update its own beliefs based on imagined beliefs of the other agents.
@@ 2 Related Works
Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) have proven effective in image generation. To render high-resolution images, the latent diffusion models (LDMs) (Rombach et al., 2022) are proposed to denoise in the latent space. Similarly, video diffusion models (Blattmann et al., 2023b; Wang et al., 2023a; Blattmann et al., 2023a; Song et al., 2025) use VAE models to encode video frames and denoise in the latent space. For controllable synthesis, the conditional denoising autoencoder are implemented with text (Rombach et al., 2022; OpenAI, 2024) and various conditioning controls (Zhang et al., 2023; Sudhakar et al., 2024). We focus on video generation conditioned on the egocentric panoramic view of agent, as panorama (Li & Bansal, 2023; 2024) ensures coherence in the generated world, and the use of egocentric vision is de facto choice in many embodied tasks (Das et al., 2018; Sermanet et al., 2024; Song et al., 2025).
Generative vision for embodied decision making. Decision-making in the physical world (Das et al., 2018; Sermanet et al., 2024) is a fundamental AI challenge. LLMs provide linguistic reasoning that aids decision-making (Hao et al., 2023; Min et al., 2024) and vision-language planning (Cen et al., 2024). World models offer predictive representations of future states to inform decisions, though early attempts (Ha & Schmidhuber, 2018; LeCun, 2022) focus on simple game agents and often lack commonsense reasoning about the physical world. Generative vision (OpenAI, 2024; Kondratyuk et al., 2024) and in-context learning (Bai et al., 2024; Zhang et al., 2024) offer new avenues for using video generation to guide real-world decision-making (Yang et al., 2024c). Several works focus on specific application domains such as autonomous driving (Hu et al., 2023; Wang et al., 2023b; 2024c; Gao et al., 2024a; b) which limit their generality. Others like video in-context learning (Zhang et al., 2024) requires a known demonstration video, which is inefficient for decision-making. Action-conditioned video generation models (Du et al., 2024a; Yang et al., 2024b; c; Wang et al., 2024a; Bu et al., 2024; Souček et al., 2024; Du et al., 2024b) can directly synthesize visual plans for decision-making. These models, however, focus on state transition probabilities without explicitly modeling agent beliefs, which are crucial for reasoning about other objects/agents in partially observable environments.
@@ 3 Generative World Exploration
Figure 2: GenEx is able to explore an imaginative world by generating video sequence, given RGB observations, exploration direction, and distance (a). GenEx, grounded in physical environment, can perform GPT-assisted goal-agnostic imaginative exploration of the world (b) and goal-driven imaginative exploration (c).
@@ 3.1 Macro-Design of GenEx
(a) Overview. As shown in  Fig. 2, the GenEx framework enables agents to explore within an imaginative world by streaming video generation, based on current RGB observations and given exploration configurations. The RGB observation is represented as a panorama image sampled from any location in the world. A large multimodal model (LMM) serves as the pilot, or the decision maker, to set up exploration configurations, including any $360^{\circ}$ navigation direction and distance. GenEx processes the input in two steps. First, it takes the exploration orientation to update the panorama forward view. Secondly, its built-in diffuser generates the forward navigation video. Both view update and diffuser are detailed in § 3.2.
@@ 3.2 Micro-Design of GenEx
We detail our micro-design as follows. Our diffuser builds upon the standard stable video diffusion (SVD) (Blattmann et al., 2023a) backbone (a), but adapts the input from conventional images to panoramas (b). Furthermore, we propose spherical-consistent learning (c) to ensure coherence in imaginative exploration.
(a) Diffuser backbone. To support exploration illustrated in Fig. 2, we propose a video diffuser that can be seamlessly adapted to be a world explorer. Given an initial panorama image $x^{0}$ with camera position $p^{0}$, our objective is to generate a sequence of images $\{x^{1},\dots,x^{n}\}$ corresponding to a sequence of camera positions $\{p^{1},\dots,p^{n}\}$. The camera positions progress steadily forward, representing navigation in the world. Since the panorama image represents a 360-degree view, the generation should persist the information stored in previous frames to maintain world consistency throughout the sequence. Our model uses the pretrained SVD (Blattmann et al., 2023a). The Transformer UNet (Ronneberger et al., 2015; Chen et al., 2021) architecture is as described in Blattmann et al. (2023b), where temporal convolution and attention layers are inserted after every spatial convolution and attention layer. The pipeline of our model is shown in Fig. 3 (a). Given an image condition $c$ (encoded from image $x^{0}$ using CLIP image Transformer (Radford et al., 2021)), the video diffusion algorithms learn a network $\epsilon_{\theta}$ to predict the noise added to the noisy image latent $z_{t}$ with $\mathcal{L}_{\text{noise}}=\|\epsilon_{\theta}(z_{t},c)-\epsilon_{t}\|^{2}$.
(b) Input image representation. Panorama images are well-suited for generative exploration as they captures all perspectives from an egocentric viewpoint into a 2D image. Essentially, it represents a spherical polar coordinate system $\mathcal{S}$ on a 2D grid in the Cartesian coordinate system $\mathcal{P}$, as shown in Fig. 3 (b). Panorama effectively stores every perspective of the world from a single location which preserves the global context during spatial navigation. This allows us to maintain consistency in world information from the conditional image, ensuring that the generated content aligns coherently with the surrounding environment. The panorama image also allows for rotational transformations, which facilitate world navigation by enabling us to rotate the image to face a different angle while preserving its original information. The rotation can be performed using Eq. 1:
@@
where $u$ and $v$ are positions on the 2D image plane, and $\phi$ and $\theta$ represent longitude and latitude in polar coordinates. The rotation function $\mathcal{R}$ applies a rotation to the spherical representation in any direction, simulating turning around during navigation. Additionally, a panorama image can be converted into a cubemap of six separate regular images, each representing a face of a cube (front, back, left, right, top, and bottom). This panorama-to-cube transformation enhances visual understanding by LMM agents. Full mathematical details of the equirectangular projection are in § A.1.
@@
[…]
@@ 4.1 Imagination-driven Belief Revision
Embodied agents operate under a POMDP framework (Puterman, 1994; Kaelbling et al., 1998). At each time step $t$, the agent’s world state (which represents the complete environment at this specific moment), $s^{t}\in S$, and action $a^{t}\in A$ determine the next world state via the transition probability ${T}(s^{t+1}|s^{t},a^{t})$. The agent’s given goal $g\in G$ (e.g., crossing the street) influences the reward $r^{t}=R(s^{t},a^{t},g)$, which drives the agent to achieve its objective. The agent receives an observation $o^{t}\in\Omega$ based on the observation model $O(o|s^{t})$ and maintains a belief, represented by a distribution $b(s)$, which is the agent’s internal estimate of the true state of the world. Its belief is updated with new observations, following the POMDP framework in Eq. 3:
$b^{t+M}(s^{t+M})=\prod_{t}^{M}\underbrace{O(o^{t+1}|s^{t+1},a^{t})\sum_{s^{t}}T(s^{t+1}|s^{t},a^{t})}_{\text{Physical Exploration}}b^{t}(s^{t})$ (3)
The decision $a^{t}$ made at any time $t$ becomes more informed as the agent gains a clearer understanding of its surroundings. By navigating through physical space, the agent gathers additional information about its environment (Fan et al., 2024), enabling more accurate assessments and better choices moving forward. However, physically traversing the space is inefficient, expensive, and even impossible in dangerous scenario. To streamline this process, we can use imagination as a medium for the agent to simulate outcomes without physically traversing. The key question becomes:
@@
Imagination-driven belief revision. We propose imagination-driven belief revision that uses imaginative exploration to enhance POMDP agents with a instant belief revision between time steps. In imagination, we freeze the time and create an imagined world, thus dropping the time variable $t$ and defining an imagination space with hat $\hat{}$ on the variables. Here the agent can make a sequence of imaginative actions $\hat{\mathbf{a}}=\{\hat{a}_{i}\in\hat{A}\}$ over imagination time step $I=\{1,...i,...,n\}$. The agents can make sequential speculation on the unobserved world based on its initial belief and toward ultimate goal, it imagine novel observations in a previous unobserved world with $p_{\theta}(\hat{o}^{i+1}|\hat{o}^{i},{\hat{a}^{i}})$, where $\hat{o}^{0}=o^{0}$ as initialization and $\theta$ is our parameterized video diffuser generating the imaginative observations. As a result, it can update its belief with Eq. 4:
$\hat{b}^{t}(s^{t})=\prod_{i}^{I}\underbrace{p_{\theta}(\hat{o}^{i+1}|\hat{o}^{i},\hat{a}^{i})}_{\text{Imaginative Exploration}}b^{t}(s^{t})$ (4)
@@
The agent make actions based on its belief and goal, with policy model $\pi(a^{t}|b^{t}(s^{t}),g)$. Through the revised belief, the agent is capable of making more informed decision toward $a^{*}$ with a more refined belief toward $b^{*}$, with more information on the true state of its surrounding environment.
In our work, we apply GenEx for imaginative exploration and a LMM as the policy model $\pi$ and belief updater $b(s)$, mapping observation to belief, with examples in Fig. 5 and system pipeline in § A.5.3.
@@ 4.2 Generalized to Multi-agent
Thus, we can adjust agent-$1$’s beliefs by aggregating the imagined belief counterpart for other K1K-1 agents.
@@ 5.1 Dataset construction
GenEx-DB. We synthesize a large-scale dataset generated using Unity, Blender, and Unreal Engine. The full details are in § A.3. We create four distinct scenes, each representing a different visual style (Realistic, Animated, Low-Texture, and Geometry), shown in Fig. 6: We train a model with each dataset, and for the four resulting navigational video diffusers, we conduct cross-validation across all scenes to evaluate their generalization capabilities (detailed in § A.7).
@@
GenEx-EQA. Through the proposed GenEx model, the agents perform exploration autonomously, capable of tackling embodied tasks such as single agent scene understanding and multi-agent intersectional reasoning. Following the four conditions in § 4.3, we design over 200 scenarios in virtual physical engine to test various LMM agents embodied decision-making. We provide comprehensive details in § A.4.1. The dataset generally represent two scenarios:
@@ 5.2 Evaluation on Generation Quality
We adopt FVD (Unterthiner et al., 2019), SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018), and PSNR (Horé & Ziou, 2010) to evaluate video generation quality, with details in § A.5.
@@
As a strong baseline, we develop a six-view navigator by training six separate diffusion model for each face of the cube, representing still a 360 view, independently (See Fig. 3 (b) six-view). The implementation detail is shown in § A.6. This baseline may align well with 2D diffusion models but stands in contrast to the panoramic approach, which is particularly effective at maintaining consistent environmental context. To enable a fair comparison with GenEx in video quality evaluation, the six-view baseline predictions are reprojected into panoramas. As a result, Table 1 shows that our method achieves high generation quality and surpass six-view baseline in all metrics.
Figure 7: Example randomly sampled trajectory for loop consistency. It forms a closed loop within the scene, with 9 rotations and in 15 meters distance.